A weight lifter pushes dumbbells of total mass 20 kg upward with a force of just a little more than the weight of the dumbbells a distance of 1.2 meters. He then lets the dumbbells drop back down without doing any work on them. If he wants to use (actually:convert) 50 kCal of energy in this workout, how many repetitions does he have to do? (There are 4.2 Joules in a calorie and 1000 calories in a kCal)
Answer #1

work done = mgh = 20 x 9.8 x 1.2 =235.2

work (energy ) needed = 50 x 10^3 x 4.2 J = 210000 J
repetitions = 210000 / 235.2 = 892.86